I need help solving a substitution problom
1/2x+y=2
2x+3y=9

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

solving a substitution problom
(1/2)x+y=2
2x+3y=9
---------------
y = 2-(1/2)x
Substitute for "y" in the 2nd foundation:
----------------
2x + 3(2-(1/2) = 9
2x + 6 - (3/2)x = 9
(1/2)x = 3
x = 6
--------
Since y = 2-(1/2)x, y = 2-(1/2)6 = -1
===============
Ans: x=6, y= -1
===================
Cheers,
Stan H.

Answer by Alan3354(69443)

I need help solving a substitution problom
1/2x+y=2
2x+3y=9
-----------
1/2x+y=2
x + 2y = 4
x = 4-2y
Sub into 2nd eqn
2*(4-2y) + 3y = 9
8-4y + 3y = 9
8 - y = 9
y = -1
---------
2x+3y = 9
2x - 3 = 9
2x = 12
x = 6